A circular coil of radius 2.0 cm has 500 turns in it and carries a current of 1.0 A. Its axis makes an angle of 30° with the uniform magnetic field of magnitude 0.40 T that exists in the space. Find the torque acting on the coil.

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. n = 500, r = 0.02 m, θ = 30° i = 1A, B = 4 × 10^–1 T i = μ × B = μ B Sin 30° = ni AB Sin 30° = 500 × 1 × 3.14 × 4 × 10^–4 × 4 × 10^–1 × (1/2) = 12.56 × 10^–2 = 0.1256 ≈ 0.13 N-M